OpenStudy (anonymous):

Algebra 2 word problem? multiple choice! Two runners leave from the same location traveling in opposite directions. The first runner is running at 7 mph and the second is running at 9 mph. How long have they been running when they are 24 miles apart? 1 hour 1.5 hours 2 hours 2.5 hours

OpenStudy (anonymous):

this is for a quiz..?

OpenStudy (anonymous):

relative speed = 16mph , relative distance=24 mils get time now using dist = speed X time

OpenStudy (ganpat):

speed = distance /time.. relative speed = (9 + 7) = 16 miles/hr.. So 16 = 24/t t = 24/16 = 1.5 hrs...

hartnn (hartnn):

take half hour window...in half hour they will be (7+9)/2=8 miles apart,then in next half hour,they will be other 8 miles apart...so on.,they will be 24 miles away in 3 half hours.....